hi
really need help...thanks a bunch
In five years annie's age will be one half of Sean's age. IN twelve years, one half of Annie's age will be one third of Sean's age. How old are they now?
thanks a lot

Answer by ptaylor(2198)

Let x=annie's age now
And let y=Sean's age now
Now we are told that:
x+5=(1/2)(y+5) or
x+5=y/2 + 5/2 multiply each side by 2
2x+10=y+5 subtract 10 and also y from each side
2x-y+10-10=y-y-10+5 collect like terms
2x-y=-5------------------------------------------------eq1
We are also told that:
(1/2)(x+12)=(1/3)(y+12)
x/2 + 6=y/3 + 4 multiply each side by 6
3x+36=2y+24 subtract 2y and also 36 from each side
3x+36-36-2y=2y-2y-36+24 collect like terms
3x-2y=-12----------------------------------------------eq2
Multiply eq1 by 2:
4x-2y=-10-------------eq1a
subtract this from eq2
-x=-2 divide each side by -1
x=2 years old-----------------annie's age now
substitute x=2 into eq2
3*2-2y=-12
6-2y=-12 subtract 6 from each side
6-6-2y=-12-6 collect like terms
-2y=-18 divide both sides by -2
y=9 years old ------------------------Sean's age now

CK
2+5=(1/2)(9+5)
7=7
also
(1/2)(2+12)=(1/3(9+12)
7=7

Hope this helps----ptaylor